OpenStudy (anonymous):

solve the following DE by using the appropriate substitution: (y^2 + yx) dx - x^2 dy = 0

OpenStudy (michele_laino):

hint: I can rewrite your equation as below: \[\Large \frac{{dy}}{{dx}} = \frac{y}{x} + {\left( {\frac{y}{x}} \right)^2}\]

OpenStudy (michele_laino):

@MarcLeclair

OpenStudy (michele_laino):

so, please try this substitution: \[\Large \frac{y}{x} = u\left( x \right)\] where u(x) is a new function of the variable x

OpenStudy (anonymous):

my question was centered more towards how do i know the substitution is possible. like M(tx,ty) = t m(x,y) but in this case the M doesnt have the same degree

OpenStudy (irishboy123):

you know this is homogeneous from the test you provided (y^2 + yx) dx - x^2 dy = 0 dy/dx = (y^2 + yx)/x^2 = f(x,y) f(kx,ky) = (k^2y^2 + kykx)/k^2x^2 = (y^2 + yx)/x^2 = f(x,y) you can also tell by inspection as the polynomials, ignoring the y' [or dy, dx], are of order 2. so v = y/x is the sub to use.